OpenStudy (anonymous):

how do I put this into scientific notation: -1456.90

OpenStudy (ash2326):

-1456.90 -1.45690* 10^3

OpenStudy (anonymous):

-1.45690x 10^3

OpenStudy (unklerhaukus):

\[-1.4569 \times 10^3\]

OpenStudy (anonymous):

but doesn't the coefficient have to be \[\le1\] and \[\ge10\]?

OpenStudy (ash2326):

no, there's no condition like that

OpenStudy (anonymous):

thats what I have in my notes...

OpenStudy (ash2326):

what's written?

OpenStudy (anonymous):

what I said is what is written

OpenStudy (anonymous):

oh, but the signs are supposed to be switched

OpenStudy (anonymous):

its\[\ge1\] and \[\le 10\]

OpenStudy (ash2326):

a* 10^b see the coefficient should be in the range \[1 \le a <10\] a minus can come before it if it's a negative number here the coefficient is 1.4569 with a negative sign

OpenStudy (anonymous):

ok thanks

OpenStudy (ash2326):

did you get it??

OpenStudy (anonymous):

yep